A magnetic field exerts a force of 8.0 x 10^–14 N toward the west on a proton moving vertically upward at a speed of 5.0 x 10^6 m/s. When moving straight in a northerly direction, the force on the proton is zero. Determine the magnitude and direction of the magnetic field B in this region. (The charge of a proton = 1.6 x 10-19 C, the north direction is into the page.).

1
Expert's answer
2022-02-22T04:41:28-0500

The magnetic force

"F=qvB"

Hence, the magnitude of the magnetic field

"B=\frac{F}{qv}\\
=\frac{8.0*10^{–14}}{1.6*10^{–19}*5.0*10^{6}}=0.1\:\rm T"

The field is directed into the page.